OpenStudy (anonymous):

The base of a triangleis 6cm more than the height. If the area of the triangle is 140 square cm, what is the length of the base?

OpenStudy (anonymous):

h = height b = h+6 140 = (1/2)(h^2 + 6h) h = 14 b = 20 The base is 20 cm